Let f (x)=|x-4| and g(x)+ 2.Find f o g and g o f
Thank you for help

Answer by stanbon(57290)

Let f (x)=|x-4| and g(x)= 2.Find f o g and g o f
--------------
f[g(x)] = f[2] = abs(2-4) = 2
==============
Stan H.